OpenStudy (anonymous):

Find the distance between points P(8, 2) and Q(3, 8) to the nearest tenth

OpenStudy (ivettef365):

do you have the distance formula

OpenStudy (anonymous):

so i just use distance formula?

OpenStudy (ivettef365):

right, on that website you enter the points and it will calculate the distance

OpenStudy (anonymous):

..or just do (8-3)^2 + (2-8)^2 ... = 25 + 36 = 61 sqrt(61) is roughly 7.8 id say